# Unified Kraft Break at $\sim\,$6500 K: A Newly Identified Single-Star Obliquity Transition Matches the Classical Rotation Break

### v1

@@ 6 Conclusion
It is a common working assumption that the stellar-obliquity transition coincides with the rotational Kraft break, motivating the view that obliquity damping in cool stars drives alignment (Winn et al., 2010a; Albrecht et al., 2012; Winn & Fabrycky, 2015; Albrecht et al., 2022). However, it has been noted (Beyer & White, 2024), and our analysis confirms, that the commonly quoted obliquity break near $T_{\rm eff}=6100~\mathrm{K}$ is several hundred kelvin cooler than the classical rotational break around $6500~\mathrm{K}$, posing a fundamental inconsistency.
@@
    A cooler break arises in multiples. The $\sim 6100~\mathrm{K}$ obliquity break appears when binaries/multiples are included (panels a and d of Figure 1), although the reason is ambiguous (See section 4 for a detailed discussion).
@@
We thank the anonymous reviewer for their constructive feedback, which has greatly improved the quality of this manuscript. We thank Jie Yu and Zheng Guo for bringing the Beyer & White (2024) paper to our attention. We are grateful for helpful discussions with Constantine Deliyannis, Caty Pilachowski, Meng Sun, Cristobal Petrovich, Brandon Radzom, Gongjie Li, Malena Rice, Samir Salim, Zachary Maas, and Timothy Bedding. This work was supported in part by the NASA Exoplanets Research Program NNH23ZDA001N-XRP (Grant No. 80NSSC24K0153), the NASA TESS General Investigator Program, Cycle 7, NNNH23ZDA001N-TESS (Grant No. 80NSSC25K7912), and the Heising-Simons Foundation (Grant #2023-4050). Additionally, X.Y.W. acknowledges support from the Sullivan Prize Fellowship. J.M.J.O. acknowledges support from NASA through the NASA Hubble Fellowship grant HST-HF2-51517.001, awarded by STScI, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. This research was supported in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute.
Table 1: Misaligned Hot Jupiters in Binaries or Multi-star systems footnotetext: Benomar et al. (2014), 2 Albrecht et al. (2012), 3 Rice et al. (2024), 4 This work, 5 Torres et al. (2010b), 6 Winn et al. (2011), 7 Ngo et al. (2015), 8 Mortier et al. (2013), 9 Johnson et al. (2011), 10 Zhao et al. (2014), 11 Eastman et al. (2016), 12 Knudstrup et al. (2024), 13 McLeod et al. (2017), 14 Rubenzahl et al. (2024), 15 Johns et al. (2019), 16 Giacalone et al. (2025), 17 Leonardi et al. (2024), 18 Bourrier et al. (2024), 19 Ngo et al. (2016), 20 Fu et al. (2021), 21 Stassun et al. (2017), 22 Neveu-VanMalle et al. (2014), 23 Jansen & Kipping (2020), 24 Addison et al. (2018), 25 Lam et al. (2017), 26 El-Badry et al. (2021), 27 Temple et al. (2019), 28 Hartman et al. (2015), 29 Smith et al. (2019).
@@
b Given that Eastman et al. (2016) estimated KELT-4 B and C to be twin K-type stars, we assume they share identical properties.
c This value was derived using lofti_\_gaia (Pearce et al., 2020), incorporating the companion mass reported by Rice et al. (2024) and astrometric measurements (relative RA and DEC, relative proper motions, and Gaia parallax) from the Gaia DR3 catalog (Gaia Collaboration et al., 2023).
@@ Appendix A Multi-component SED modeling for systems with stellar companions
For HAT-P-7, HAT-P-14, HAT-P-30, WASP-12, WASP-100, and WASP-136, we performed multi-component SED modeling using EXOFASTv2 (Eastman, 2017; Eastman et al., 2019) to derived non-contamination SED-based $T_{\rm eff}$  for primary stars. We adopted broadband photometry including Gaia $G$, $G_{\rm bp}$, $G_{\rm rp}$, 2MASS $JHK$, and WISE 1–3 for primary stars. For the stellar companions, we included broad photometry from Gaia DR3 together with direct imaging measurements (Ngo et al., 2015; McLeod et al., 2017). Gaussian priors on $\left[{\rm Fe}/{\rm H}\right]$ were derived from the SWEET-Cat catalog (Santos et al., 2013; Sousa et al., 2021), which provides spectroscopic stellar parameters derived from a uniform and homogeneous analysis of the spectra. We assumed that the primary and companion stars share the same initial $\left[{\rm Fe}/{\rm H}\right]$, $V$-band extinction, distance, and age, and adopted the $V$-band extinction value from Schlafly & Finkbeiner (2011) as the upper limit on extinction. The Gaia DR3 parallax corrected for the zero-point offset, along with its uncertainty (Gaia Collaboration et al., 2023; Lindegren et al., 2021), was adopted as a Gaussian prior on the parallax.
WASP-180 A and B are resolved in Gaia DR3, 2MASS, and WISE, but no $T_{\rm eff}$  from SED solution is available. We therefore performed single-component SED fit using the aforementioned broad photometry. For KELT-18, although McLeod et al. (2017) performed multi-component SED modeling, they did not report the stellar companion’s mass. We therefore carried out our own multi-component SED modeling to derive it. K2-237 lies in a crowded stellar field, making it impossible to identify bound companions. Therefore, we treated it as a single star and performed single-star SED modeling using Gaia $G$, $G_{\rm bp}$, $G_{\rm rp}$, and 2MASS $JHK$, excluding WISE 1–3 photometry due to potential light contamination from nearby stars.
We adopted standard convergence criteria (independent draws $T_{z}$ >> 1000 and Gelman–Rubin statistic $\hat{R}$< 1.01; Eastman et al. 2019). Our analysis, summarized in Table 1, shows that, for primary stars, the $T_{\rm eff}$ values from the uncontaminated SED modeling remain consistent, within $3\sigma$, with the $T_{\rm eff}$  values adopted in this work from TEPCat.

### v2

@@ VI Conclusion
It is a common working assumption that the stellar-obliquity transition coincides with the rotational Kraft break, motivating the view that obliquity damping in cool stars drives alignment (Winn2010; Albrecht2012; WinnFabrycky2015; Albrecht2022). However, it has been noted (Beyer2024), and our analysis confirms, that the commonly quoted obliquity break near $T_{\rm eff}=6100~\mathrm{K}$ is several hundred kelvin cooler than the classical rotational break around $6500~\mathrm{K}$, posing a fundamental inconsistency.
@@
    A cooler break arises in multiples. The $\sim 6100~\mathrm{K}$ obliquity break appears when binaries/multiples are included (panels a and d of Figure 1), although the reason is ambiguous (See section IV for a detailed discussion).
@@
We thank the anonymous reviewer for their constructive feedback, which has greatly improved the quality of this manuscript. We thank Jie Yu and Zhao Guo for bringing the Beyer2024 paper to our attention. We are grateful for helpful discussions with J. J. Zanazzi, Eugene Chiang, Yubo Su, Yanqin Wu, Janosz Dewberry, Fred C. Adams, Simon H. Albrecht, Joshua N. Winn, Fei Dai, Constantine Deliyannis, Caty Pilachowski, Kareem El-Badry, Cheyanne Shariat, Gang Li, Meng Sun, Cristobal Petrovich, Brandon Radzom, Gongjie Li, Malena Rice, Samir Salim, Zachary Maas, and Timothy Bedding. This work was supported in part by the NASA Exoplanets Research Program NNH23ZDA001N-XRP (Grant No. 80NSSC24K0153), the NASA TESS General Investigator Program, Cycle 7, NNNH23ZDA001N-TESS (Grant No. 80NSSC25K7912), and the Heising-Simons Foundation (Grant #2023-4050). S.W. gratefully acknowledges support from the John and A-Lan Reynolds Faculty Research Fund, which enabled participation in the International Conference on Exoplanets and Planet Formation (EPF). This support was vital for presenting this result and discussing the puzzling observation with the world’s leading researchers in the field. Additionally, X.Y.W. acknowledges support from the Sullivan Prize Fellowship. J.M.J.O. acknowledges support from NASA through the NASA Hubble Fellowship grant HST-HF2-51517.001, awarded by STScI, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. This research was supported in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute.
Table 1: Misaligned Hot Jupiters in Binaries or Multi-star systems footnotetext: Benomar2014, 2 Albrecht2012, 3 Rice2024, 4 This work, 5 Torres2010HATP14, 6 Winn2011, 7 Ngo2015, 8 Mortier2013, 9 Johnson2011, 10 Zhao2014, 11 Eastman2016, 12 Knudstrup2024, 13 McLeod2017, 14 Rubenzahl2024, 15 Johns2019, 16 Giacalone2025, 17 Leonardi2024, 18 Bourrier2024, 19 Ngo2016, 20 Fu2021, 21 Stassun2017, 22 NeveuVanMalle2014, 23 Jansen2020, 24 Addison2018, 25 Lam2017, 26 Badry2021, 27 Temple2019, 28 Hartman2015, 29 Smith2019.
@@
b Given that Eastman2016 estimated KELT-4 B and C to be twin K-type stars, we assume they share identical properties.
c This value was derived using lofti_\_gaia (Pearce2020), incorporating the companion mass reported by Rice2024 and astrometric measurements (relative RA and DEC, relative proper motions, and Gaia parallax) from the Gaia DR3 catalog (GaiaCollaboration2023).
@@ Appendix A Multi-component SED modeling for systems with stellar companions
For HAT-P-7, HAT-P-14, HAT-P-30, WASP-12, WASP-100, and WASP-136, we performed multi-component SED modeling using EXOFASTv2 (Eastman2017; Eastman2019) to derived non-contamination SED-based $T_{\rm eff}$  for primary stars. We adopted broadband photometry including Gaia $G$, $G_{\rm bp}$, $G_{\rm rp}$, 2MASS $JHK$, and WISE 1–3 for primary stars. For the stellar companions, we included broad photometry from Gaia DR3 together with direct imaging measurements (Ngo2015; McLeod2017). Gaussian priors on $\left[{\rm Fe}/{\rm H}\right]$ were derived from the SWEET-Cat catalog (Santos2013sweetcat; Sousa2021), which provides spectroscopic stellar parameters derived from a uniform and homogeneous analysis of the spectra. We assumed that the primary and companion stars share the same initial $\left[{\rm Fe}/{\rm H}\right]$, $V$-band extinction, distance, and age, and adopted the $V$-band extinction value from Schlafly2011 as the upper limit on extinction. The Gaia DR3 parallax corrected for the zero-point offset, along with its uncertainty (GaiaCollaboration2023; Lindegren2021), was adopted as a Gaussian prior on the parallax.
WASP-180 A and B are resolved in Gaia DR3, 2MASS, and WISE, but no $T_{\rm eff}$  from SED solution is available. We therefore performed single-component SED fit using the aforementioned broad photometry. For KELT-18, although McLeod2017 performed multi-component SED modeling, they did not report the stellar companion’s mass. We therefore carried out our own multi-component SED modeling to derive it. K2-237 lies in a crowded stellar field, making it impossible to identify bound companions. Therefore, we treated it as a single star and performed single-star SED modeling using Gaia $G$, $G_{\rm bp}$, $G_{\rm rp}$, and 2MASS $JHK$, excluding WISE 1–3 photometry due to potential light contamination from nearby stars.
We adopted standard convergence criteria (independent draws $T_{z}$ >> 1000 and Gelman–Rubin statistic $\hat{R}$< 1.01; Eastman2019). Our analysis, summarized in Table 1, shows that, for primary stars, the $T_{\rm eff}$ values from the uncontaminated SED modeling remain consistent, within $3\sigma$, with the $T_{\rm eff}$  values adopted in this work from TEPCat.